A trolley of mass 5 kg on a horizontal smooth surface is pulled by a load of mass 2 kg by means of uniform rope ABC of length 2m and mass 1kg  .As the load falls from BC=0 to BC=2m .its acceleration in  m/s changes
 

Grade:11

1 Answers

Arun
3 years ago
Initial force = 2 g = 20 N
 
Total mass = 5 +  = 6 Kg
 
BC = 2 m
 
Initial acceleration = Force Mass=20 / 5 + 1=20 / 6 ms−2
 
Final force = (Load + mass of thread)×g
 
                  = (2+1) ×10=30 N
 
Therefore, final acceleration =30 / 6 ms^−2.
 
Thus the acceleration changes from 30 / 6 ms^−2.